Please help me with this problem: (-5+x)/8 = 2x. Help would be greatly appreciated. Thanks
Answer by ilana(307)

We can begin by multiplying both sides by 8 (or "cross-multiplying"), giving -5+x=16x. I like to rewrite this as x-5=16x. Now, we have x on both sides, so we isolate x by subtracting it from both sides, giving -5=15x, and then dividing by 15, giving -5/15=x, or x=-1/3.
We can check this solution in the original:
(-5+(-1/3))/8=2(-1/3)
(-16/3)/8=-2/3
(-16/3)*(1/8)=-2/3
-16/24=-2/3
-2/3=-2/3, so this works!:)